OpenStudy (anonymous):

Factor: 4x^2-12x+9=0

OpenStudy (anonymous):

x = 1.5

OpenStudy (saifoo.khan):

wrong^

OpenStudy (anonymous):

discriminant is 0 so 1 answer

hero (hero):

4x^2 - 6x-6x + 9 = 0 2x(2x-3)-3(2x-3) = 0 (2x-3)(2x-3) = 0 x = 3/2

OpenStudy (saifoo.khan):

Right^

hero (hero):

lol x = 3/2 = 1.5

OpenStudy (anonymous):

ahahahah =.=

hero (hero):

So Jessica was right

OpenStudy (saifoo.khan):

lol, i know she's right obv. i was jk. :P

hero (hero):

I'll give you a pass on that one, but next time.....

OpenStudy (saifoo.khan):

....?

OpenStudy (anonymous):

next time you get a spank :P

hero (hero):

When it happens, you'll see

OpenStudy (saifoo.khan):

what if i don't see? what if i close my eyes?

hero (hero):

You can't keep your eyes closed forever

OpenStudy (saifoo.khan):

i never said i will close my eyes forever. :p

hero (hero):

You can close your eyes, but when you open them, you will see

hero (hero):

That is a true statement...irrefutable

OpenStudy (saifoo.khan):

lol, you will actually bothered waiting for soooo ling.

OpenStudy (saifoo.khan):

long*

OpenStudy (saifoo.khan):

Prove it that it is true.

hero (hero):

let x = eyes closed let y = eyes opened whenever x is false, y is true

OpenStudy (saifoo.khan):

what about the z?

OpenStudy (saifoo.khan):

One eye close one open?

hero (hero):

z is undefined

hero (hero):

because I didn't define it

OpenStudy (saifoo.khan):

im defining it, so it isnt.

hero (hero):

You can't change my proof. Because it's mine. Your eyes, but my proof

OpenStudy (saifoo.khan):

that dosn't makes any sense. sorry for that. :p

hero (hero):

It makes sense to me. That's all that matters. I don't care if it makes sense to you. It's like trying to interpret someone else's notes. Make your own.

OpenStudy (saifoo.khan):

It does not makes sense to me. That's all that matters.

hero (hero):

If something doesn't make sense to you, you should go find a tutor.

OpenStudy (saifoo.khan):

lol, you should see find a doctor to get your brain checked.

hero (hero):

You first

OpenStudy (saifoo.khan):

No you first.

OpenStudy (saifoo.khan):

Elders first.

hero (hero):

I'm not an Elder

hero (hero):

You can't steal my joke. I said you first. Think of something else

OpenStudy (saifoo.khan):

that is not even a joke.

OpenStudy (saifoo.khan):

Yes u r elder than me

hero (hero):

the correct word is "older"

hero (hero):

But not really, because the word old isn't associated with me. I'm immortal

OpenStudy (saifoo.khan):

elder is also a word. such a dumb.

hero (hero):

It doesn't hurt to say something nice every once in a while. I know it would be way out of your way to say something nice to someone other than girls.

hero (hero):

This proves that you're a hater